OpenStudy (anonymous):

The bank has accepted Mrs. Hudson's application for a mortgage to buy property worth $375,575. She has to pay 19 percent of the purchase price in cash before closing the deal. Calculate the loan amount the bank will give her. A. $302,415.75 B. $304,215.75 C. $340,215.75

OpenStudy (anonymous):

0.19*375,575

OpenStudy (anonymous):

X % of Y= \[\frac{ X }{ 100 }* Y\]

OpenStudy (anonymous):

So I Will Multiply 0.19 By 375

OpenStudy (anonymous):

no 0.19 by 375,575

OpenStudy (anonymous):

So Am I Dividing

OpenStudy (anonymous):

no first do 0.19* 375,575 and then subtract it from 375,575

OpenStudy (anonymous):

Can You Do It I Dont Understand

OpenStudy (anonymous):

0.19*375,575= 71359.25 then loan amount = 375,575-71359.25=

OpenStudy (anonymous):

@DamnImBad does it look like B

OpenStudy (anonymous):

Yes I Got It Thannks ,

OpenStudy (anonymous):

:) YW